OpenStudy (anonymous):

Simplify i^16 A)-1 B)1 C)i D)-i

OpenStudy (astrophysics):

Hint: \[i^2 = -1\]

OpenStudy (anonymous):

X it by 16?

OpenStudy (astrophysics):

No, what if I write it as \[i^{16} = i^2 \times i^2 \times i^2....\] 8 times

OpenStudy (jhannybean):

\[(i^2)^8 =~?\]

OpenStudy (astrophysics):

\[(i^4)^4\]

OpenStudy (anonymous):

1?

OpenStudy (astrophysics):

Yup!

OpenStudy (jhannybean):

\[(i^2)^{32/4}\]

OpenStudy (jhannybean):

Lol ok im done.

OpenStudy (astrophysics):

lol

OpenStudy (anonymous):

jhannybean don't know what you meant on that last comment you put lol

OpenStudy (jhannybean):

Oh, i said im done trying to find derivations of the same thing : \(i^{16}\) lol.

OpenStudy (anonymous):

I was talking about where you got 32/4

OpenStudy (anonymous):

any exponent of i divisible by 4=1 ie i^16, i^32, i^400 all = 1

OpenStudy (anonymous):

I would had never known that if you did not just tell me so thank you:)

OpenStudy (anonymous):

:)

OpenStudy (astrophysics):

32/4 = 8